OpenStudy (anonymous):

a lot is 60 m by 32m a house 32m by 8m is built on the lot. how much area is left over

HanAkoSolo (jamierox4ev3r):

are the lots rectangles?. if so, just multiply 60*32 and 32*8

HanAkoSolo (jamierox4ev3r):

then subtract the bigger number by the smaller number.

HanAkoSolo (jamierox4ev3r):

if you do this, what is your answer @tuntun123 ?

OpenStudy (anonymous):

1664?

HanAkoSolo (jamierox4ev3r):

no, 60*32=1380, and 32*8=256 1380-256=1124 \(m^2\)

OpenStudy (anonymous):

60 times 32 is 1920

HanAkoSolo (jamierox4ev3r):

it is! lol sorry my mental math is not working today. so in that case, if you subtract 1920 and 256, you would indeed get 1664 \(m^2\)

HanAkoSolo (jamierox4ev3r):

nice job. Do you understand the concept now @tuntun123 ?

OpenStudy (anonymous):

thanks so much u up to help me with another problem

HanAkoSolo (jamierox4ev3r):

sure

HanAkoSolo (jamierox4ev3r):

but open a new post this ones getting a bit clogged on my computer

OpenStudy (anonymous):

what is the area of a section of pavement that is 20ft wide and 60yds long,, give answer in square feet

HanAkoSolo (jamierox4ev3r):

or il do it here i guess lol

OpenStudy (anonymous):

sorry

HanAkoSolo (jamierox4ev3r):

no its fine my responses will just be a bit delayed thats all

HanAkoSolo (jamierox4ev3r):

first you should convert the yards into feet. Since 1 yard=3 feet, multiply 60 by 3 60*3=180

HanAkoSolo (jamierox4ev3r):

then, 180*20=3600 \(ft^2\)

HanAkoSolo (jamierox4ev3r):

3600 \(ft^2\) is your answer. Does that make sense @tuntun123 ?

OpenStudy (anonymous):

yes thank u

HanAkoSolo (jamierox4ev3r):

no problem, and i see you are new to OpenStudy, so..\(\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\sf\color{blue}{~~~~Welcome~to~OpenStudy!~\ddot\smile}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\)

HanAkoSolo (jamierox4ev3r):

feel free to ask questions anytime, and if you want, you can mention me in a problem like this @Jamierox4ev3r and you will be more likely to have my attention.